# MSHyper: Multi-Scale Hypergraph Transformer for Long-Range Time Series Forecasting

### v1

@@ 4.4 Objective Function
where $\widehat{\mathbf{X}}_{T+1:T+H}^{\text{O}}$ and $\mathbf{X}_{T+1:T+H}^{\text{O}}$ are ground truth and forecasting results, respectively.
5 Experiments
5.1 Experimental Setup
@@
Table 1: Multivariate long-range time series forecasting results on five real-world datasets. The input length is set as $I=168$, and the prediction length $O$ is set as 96, 168, 336, and 720 (For ETTm1, the prediction length is set as 96, 288, 672, and 720). IMP shows the improvement of MSHyper over the best baseline. The best results are bolded and the second best results are underlined.
Methods
@@
(AAAI 2021)
@@
(NeurIPS 2021)
Pyraformer
(ICLR 2021)
FEDformer
(ICML 2022)
Quatformer
(KDD 2022)
Crossformer
(ICLR 2023)
@@
IMP
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather 96 0.300 0.384 0.266 0.336 0.503* 0.518* 0.217 0.296 0.211 0.279 0.444* 0.473* 0.164 0.217 22.27% 22.22%
168 0.608 0.567 0.574 0.548 0.519 0.521 0.564 0.541 0.251* 0.295* 0.473 0.494 0.197 0.246 21.51% 16.61%
336 0.702 0.620 0.600 0.571 0.539 0.543 0.533 0.536 0.310 0.344 0.495 0.515 0.262 0.298 15.48% 13.37%
720 0.831 0.731 0.587 0.570 0.547 0.553 0.562 0.557 0.381 0.374 0.526 0.542 0.309 0.325 18.90% 13.10%
Electricity 96 0.274 0.368 0.255 0.339 0.272* 0.371* 0.183 0.297 0.197 0.308 0.185* 0.278* 0.183 0.271 0% 2.52%
168 0.368 0.424 0.299 0.387 0.452 0.455 0.263 0.361 0.205* 0.315* 0.231 0.309 0.193 0.298 5.85% 3.56%
336 0.381 0.431 0.375 0.428 0.463 0.456 0.305 0.386 0.220 0.329 0.323 0.369 0.205 0.310 6.82% 5.78%
720 0.406 0.443 0.377 0.434 0.480 0.461 0.372 0.434 0.245 0.350 0.404 0.423 0.230 0.329 6.12% 6.00%
ETTh1 96 0.865 0.713 0.449 0.459 0.701* 0.635* 0.376 0.419 0.422* 0.447* 0.396* 0.412* 0.375 0.398 0.27% 3.40%
168 0.931 0.752 0.493 0.479 0.781 0.675 0.412 0.449 0.453 0.489 0.410 0.441 0.405 0.427 1.22% 3.17%
336 1.128 0.873 0.509 0.492 0.912 0.747 0.456 0.474 0.491 0.495 0.440 0.461 0.428 0.439 2.73% 4.77%
720 1.215 0.896 0.539 0.537 0.993 0.792 0.521 0.515 0.523 0.525 0.519 0.524 0.503 0.507 3.08% 1.55%
ETTm1 96 0.678 0.614 0.502 0.476 0.520 0.504 0.366 0.412 0.375 0.398 0.320 0.373 0.315 0.365 1.56% 2.14%
288 1.056 0.786 0.604 0.522 0.729 0.657 0.398 0.433 0.408 0.424 0.404 0.427 0.394 0.419 1.01% 1.17%
672 1.192 0.926 0.607 0.530 0.980 0.678 0.455 0.464 0.472* 0.468* 0.569 0.528 0.451 0.462 0.88% 0.43%
720 1.166 0.823 0.671 0.561 0.912* 0.727* 0.543 0.490 0.499 0.502 0.751* 0.677* 0.493 0.471 1.20% 3.88%
Traffic 96 0.719 0.391 0.613 0.388 0.628* 0.354* 0.587 0.366 0.618 0.384 0.450* 0.477* 0.408 0.269 9.33% 24.01%
168 0.660 0.391 0.649 0.407 0.635 0.347 0.607 0.385 0.621* 0.399* 0.513 0.289 0.418 0.271 18.52% 6.23%
336 0.747 0.405 0.624 0.388 0.641 0.347 0.624 0.389 0.622 0.384 0.530 0.300 0.431 0.287 18.68% 4.33%
720 0.792 0.430 0.674 0.417 0.670 0.364 0.623 0.378 0.629 0.383 0.573 0.313 0.463 0.298 19.20% 4.79%
  • * indicates that some methods do not have uniform prediction lengths with other methods. To ensure a fair comparison, we utilize their official code and adjust prediction lengths.
Datasets. We conduct experiments on five commonly used long-range time series forecasting datasets, including ETT (ETTh and ETTm), Electricity, Weather, and Traffic datasets. The descriptions about the five datasets are as follows:
ETT Zhou et al. (2021): This dataset contains the oil temperature and load data collected by electricity transformers, including ETTh (ETTh1 and ETTh2) and ETTm (ETTm1 and ETTm2), which are sampled hourly and every 15 minutes, respectively.
Traffic222 http://pems.dot.ca.gov: This dataset contains the road occupancy rates of 862 sensors in San Francisco Bay Area freeways, which is sampled hourly.
Weather333 https://www.bgc-jena.mpg.de/wetter/: This dataset contains 21 meteorological measurements data from the Weather Station of the Max Planck Biogeochemistry, which is sampled every 10 minutes.
Following existing works Zhou et al. (2021); Zhang and Yan (2023); Liu et al. (2021), we split each dataset into training, validation, and testing set based on chronological order. The ratio is 6:2:2 for the ETT dataset and 7:2:1 for the others.
Experimental Settings. MSHyper is trained/tested on a single NVIDIA Geforce RTX 3090 GPU. The aggregation windows are set to 4 for ETTh and ETTm datasets and 3 for other datasets. $S$ and $H$ in MSHyper are set to 4 in all experiments. Z-Score normalization is used to normalize all datasets. MSE and MAE are used as evaluation metrics. Lower MSE and MAE results mean better performance. Adam is set as the optimizer with the initial learning rate of $\text{10}^{\text{-4}}$, and the batch size is set to 32.
Baselines. The baselines include six latest Transformer-based models: Informer Zhou et al. (2021), Autoformer Wu et al. (2021), Pyraformer Liu et al. (2021), FEDformer Zhou et al. (2022), Quatformer Chen et al. (2022), and Crossformer Zhang and Yan (2023). The source code of MSHyper is released on Anonymous GitHub [^4] https://anonymous.4open.science/r/MSHyper-A3FF.
5.2 Main Results
Multivariate Results. Table 1 shows the multivariate experimental results of MSHyper compared with baselines on five datasets. We can see that MSHyper achieves state-of-the-art (SOTA) results on all five datasets at all prediction lengths. From Table 1 we can discern the following tendencies: 1) Informer and Autoformer exhibit relatively poor predictive performance. This may come from that time series forecasting requires modeling temporal pattern interactions, while vanilla attention or simplistic decomposition techniques are insufficient in capturing such interactions. 2) By considering the temporal pattern interactions of different scales, Pyraformer, Crossformer, FEDformer, and Quatformer all achieve excellent performance. 3) Benefiting from its ability to model high-order temporal pattern interactions, MSHyper achieves the best performance in long-range time series forecasting. Specifically, MSHyper reduces prediction errors by an average of 8.73% and 7.15% over the best baseline in MSE and MAE, respectively.
Univariate Results. Table 2 summaries the results of univariate long-range time series forecasting on ETTh (ETTh1 and ETTh2) and ETTm (ETTm1 and ETTm2) datasets. We can see from Table 2 that MSHyper achieves SOTA results on all datasets at all prediction lengths. Specifically, MSHyper gives an average error reduction of 9.57% and 7.09% compared to the best baseline in MSE and MAE, respectively. The experimental results verify the effectiveness of MSHyper in univariate long-range time series forecasting.
Table 2: Univariate long-range time series forecasting results on ETT dataset. The input length is set as $I$=168, and the prediction length $O$ is set as 96, 192, 336, and 720. IMP shows the improvement of MSHyper over the best baseline. The best results are bolded and the second best results are underlined.
Methods
Informer
(AAAI 2021)
Autoformer
Pyraformer*
(ICLR 2021)
FEDformer
(ICML 2022)
Quatformer*
(KDD 2022)
Crossformer*
(ICLR 2023)
MSHyper
IMP
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.193 0.377 0.071 0.206 0.099 0.277 0.079 0.215 0.069 0.222 0.076 0.216 0.060 0.187 13.04% 9.22%
192 0.217 0.395 0.114 0.262 0.174 0.346 0.104 0.245 0.138 0.222 0.085 0.225 0.078 0.215 8.24% 3.15%
336 0.202 0.381 0.107 0.258 0.198 0.370 0.119 0.270 0.123 0.275 0.106 0.257 0.087 0.226 17.92% 12.06%
720 0.183 0.355 0.126 0.283 0.209 0.384 0.142 0.299 0.146 0.293 0.128 0.287 0.108 0.230 14.29% 18.73%
ETTh2 96 0.213 0.373 0.153 0.306 0.152 0.303 0.128 0.271 0.135 0.278 0.125 0.273 0.117 0.266 6.40% 1.85%
192 0.227 0.387 0.204 0.351 0.197 0.348 0.185 0.330 0.189 0.342 0.187 0.334 0.172 0.325 7.03% 1.52%
336 0.242 0.401 0.246 0.389 0.238 0.385 0.231 0.378 0.234 0.375 0.227 0.377 0.222 0.373 2.20% 0.53%
720 0.291 0.439 0.268 0.409 0.274 0.435 0.278 0.420 0.263 0.424 0.266 0.410 0.242 0.401 7.98% 1.96%
ETTm1 96 0.109 0.277 0.056 0.183 0.127 0.281 0.033 0.140 0.035 0.147 0.035 0.145 0.032 0.136 3.03% 2.86%
192 0.151 0.310 0.081 0.216 0.205 0.343 0.058 0.186 0.174 0.193 0.055 0.180 0.049 0.168 10.91% 6.67%
336 0.427 0.591 0.076 0.218 0.302 0.457 0.084 0.231 0.105 0.237 0.072 0.209 0.066 0.195 8.33% 6.70%
720 0.438 0.586 0.110 0.267 0.387 0.485 0.102 0.250 0.112 0.264 0.097 0.248 0.087 0.224 10.31% 9.68%
ETTm2 96 0.088 0.225 0.065 0.189 0.074 0.208 0.067 0.198 0.075 0.217 0.058 0.183 0.047 0.165 18.97% 9.84%
192 0.132 0.283 0.118 0.256 0.116 0.252 0.102 0.245 0.117 0.263 0.105 0.237 0.094 0.212 7.84% 10.55%
336 0.180 0.336 0.154 0.305 0.143 0.295 0.130 0.279 0.145 0.286 0.133 0.280 0.123 0.242 5.38% 13.26%
720 0.300 0.435 0.182 0.335 0.197 0.338 0.178 0.325 0.183 0.368 0.181 0.324 0.158 0.308 11.24% 4.94%
  • * indicates that some methods do not yield results for univariate long-range time series forecasting. To ensure a fair comparison, we utilize their official code and fine-tune key hyperparameters via grid search. Other results are from FEDformer.
5.3 Ablation Studies
Multi-Scale Hypergraph. To investigate the effectiveness of the multi-scale hypergraph, we conduct ablation studies by carefully designing the following three variants: 1) Replacing the H-HGC model with the fully-connected graph (MSHyper-FCG). 2) Replacing the H-HGC model with the pyramid graph used in Liu et al. (2021) (MSHyper-PG). 3) Connecting the input sequence with intra-scale hyperedges, and thus the multi-scale hypergraph turns to the single-scale hypergraph (MSHyper-SSH). The experimental results are shown in Table 3, from which we can observe that MSHyper performs the best in all cases, which indicates the importance of multi-scale hypergraph in modeling high-order interactions between temporal patterns of different scales.
Table 3: Results of different multi-scale hypergraph construction methods.
Methods Prediction length
168 336 720
@@ 5.3 Ablation Studies
0.405
0.427
0.428
0.439
0.503
0.507
Hyperedge Graph. To investigate the effectiveness of the hyperedge graph, we conduct ablation studies by carefully designing the following three variants: 1) Removing the association relationship of the hyperedge graph (MSHyper-w/o AR). 2) Removing the sequential relationship of the hyperedge graph (MSHyper-w/o SR). 3) Removing the hyperedge graph, i.e., without the hyperedge-hyperedge phase (MSHyper-w/o H-H). The experimental results are shown in Table 4. We can observe that MSHyper performs better than MSHyper-w/o AR and MSHyper-w/o SR, showing the effectiveness of the association relationship and sequential relationship, respectively. In addition, removing the hyperedge-hyperedge phase gets the worst forecasting results, which demonstrates the superiority of the hyperedge graph in enhancing hypergraph modeling.
Table 4: Results of different hyperedge graph construction methods.
Methods Prediction length
168 336 720

### v2

@@ IV-E Objective Function
where $\widehat{\mathbf{X}}_{T+1:T+H}^{\text{O}}$ and $\mathbf{X}_{T+1:T+H}^{\text{O}}$ are forecasting results and ground truth, respectively.
IV-F Complexity Analysis
The time complexity of MSHyper consists of three main parts. For the node-hyperedge phase, the time complexity is $\mathcal{O}(M)$, where $M$ is the number of hyperedges. For the hyperedge-hyperedge phase, the time complexity is $\mathcal{O}(M^{2})$. For the hyperedge-node phase, since $\mathbf{D}_{v}$ and $\mathbf{D}_{e}$ are diagonal matrices, and the sparse operation in torch_geometric of PyTorch, the time complexity is $\mathcal{O}(MN)$, where $N$ is the number of nodes. In practical operation, since the large aggregation window $l^{s}$ and a hyperedge can connect multiple nodes (large $H_{s}$), $M$ is smaller than $N$. As a result, the total complexity of MSHyper is bounded by $\mathcal{O}(MN)$.
V Experiments
In this section, we first present experiments to verify the performance of MSHyper on eight public time series forecasting datasets. Then, we conduct ablation studies, case studies, and parameter studies to verify the effect of different module designs and parameter choices. In addition, we provide computation cost to verify the effectiveness of MSHyper.
V-A Datasets and Settings
Datasets. We conduct experiments on eight commonly used long-range time series forecasting datasets, including ETT (ETTh and ETTm), Electricity, Weather, Flight, and Exchange-Rate datasets. Table I shows the summarized dataset statistics. The auto-correlation function (ACF) values is used to evaluate the correlation between a time series and its lagged values [22]. Higher ACF values typically indicates greater predictability. The detailed descriptions about the eight datasets are given as follows:
  • ETT [60]: This dataset contains the oil temperature and load data collected by electricity transformers, including ETTh (ETTh1 and ETTh2) and ETTm (ETTm1 and ETTm2), which are sampled hourly and every 15 minutes, respectively.
  • Flight [6]: This dataset contains changes in flight data from 7 major European airports provided by OpenSky [^2] https://opensky-network.org/, which is sampled hourly.
  • Weather [^3] https://www.bgc-jena.mpg.de/wetter/: This dataset contains 21 meteorological measurements data from the Weather Station of the Max Planck Biogeochemistry, which is sampled every 10 minutes.
  • Exchange-Rate[25]: This dataset contains the exchange rate data from 8 foreign countries, which is sampled daily.
Following existing works [6, 60, 56, 28], we split each dataset into training, validation, and testing sets based on chronological order. The ratio is 6:2:2 for the ETT dataset and 7:2:1 for the others.
Experimental Settings. MSHyper is implemented in Python with PyTorch 1.13.1 and trained/tested on a single NVIDIA Geforce RTX 3090 GPU, and the source code of MSHyper is released on GitHub  [^4] https://github.com/shangzongjiang/MSHyper. We repeat all experiments 3 times and use the mean of the metrics as the final results. Following existing works [6, 29, 54], we use instance normalization to normalize all datasets. Adam is set as the optimizer. The aggregation windows are set to 4 for ETT dataset and 3 for other datasets. $S$ and $H$ in MSHyper are set to 4 in all experiments. For other hyperparameters, we use Neural Network Intelligence (NNI) [^5] https://nni.readthedocs.io/en/latest/ toolkit to automatically search the best hyperparameters, which can greatly reduce computation cost compared to the grid search approach. The detailed search space of hyperparameters and the configurations of NNI are given in Table II.
For the long-range time series forecasting, we have two kinds of settings (i.e., multivariate settings and univariate settings), which are used to forecast all feature dimensions and the last feature dimension, respectively. Following existing works[54, 61], we use all eight datasets for multivariate long-range time series forecasting and ETT dataset for univariate long-range time series forecasting, respectively.
TABLE II: Settings of NNI.
Parameters Choise
Search space Batch size {8, 16, 32, 64, 128}
Dropout rate {0.05, 0.1, 0.5}
Dimension of model {64, 128, 256, 512}
Initial learning rate {0.0001, 0.005, 0.01, 0.1}
Configures Max trial number 120
Optimization algorithm Tree-structured Parzen Estimator
Early stopping strategy Curvefitting
Evaluation metrics. Mean Square Error (MSE) and Mean Absolute Error (MAE) are used as evaluation metrics, which are defined as follows:
$\mathrm{MSE}=\frac{1}{H}\left\|\widehat{\mathbf{X}}_{T+1:T+H}^{\text{O}}-\mathbf{X}_{T+1:T+H}^{\text{O}}\right\|^{2}$ (25)
$\mathrm{MAE}=\frac{1}{H}\left\|\widehat{\mathbf{X}}_{T+1:T+H}^{\text{O}}-\mathbf{X}_{T+1:T+H}^{\text{O}}\right\|,$
where $T$ and $H$ are the input and output lengths, $\widehat{\mathbf{X}}_{T+1:T+H}^{\text{O}}$ and $\mathbf{X}_{T+1:T+H}^{\text{O}}$ are forecasting results and ground truth, respectively. Lower MSE and MAE results mean better performance.
V-B Methods for Comparison
Baselines. We carefully choose nine well-acknowledged forecasting models as our baselines, including (1) Transformer-based methods: Crossformer [56], Pyraformer [28], Autoformer [41], and Informer[60]; (2) Graph-based methods: MSGNet [6] and MTGNN [42]; (3) Linear-based methods: DLinear [54] and TiDE [15]; and (4) TCN-based methods: TimesNet [29]. The detailed descriptions about the nine baselines are as follows:
  • MSGNet [6]: It captures inter-scale correlations at different scales by leveraging frequency domain analysis and adaptive graph convolution.
  • TimesNet [29]: It transforms the 1D input sequence into 2D tensors and uses 2D convolution kernels to capture mixed-scale pattern interactions.
  • Crossformer [56]: It combines a two-stage attention with a hierarchical encoder-decoder architecture to capture cross-time and cross-dimension interactions.
  • DLinear [54]: It decomposes the input sequence into seasonal and trend components and employs two one-layer linear layers to model each component separately.
  • TiDE [15]: It uses dense MLP-based encoder-decoder architectures to handle covariates and non-linear dependencies.
  • Pyraformer [28]: It extends the two-layer structure into multi-scale embeddings, which models the interactions between nodes of different scales through a pyramid graph.
  • Autoformer [41]: It introduces an auto-correlation mechanism to realize seasonal-trend decomposition at the level of sub-sequence.
  • Informer [60]: It obtains the dominant query by calculating the KL-divergence to reduce the computational complexity.
  • MTGNN [42]: It utilizes a graph learning module to measure inter-variable dependencies and a temporal convolution module to capture temporal pattern interactions.
V-C Main Results
Multivariate Results. Table III shows the multivariate long-range time series forecasting results of MSHyper compared with baselines on all eight datasets, from which we can discern the following tendencies:
@@ V-C Main Results
TABLE III: Multivariate long-range time series forecasting results on eight real-world datasets. The input length is set as $I=96$, and the prediction length $O$ is set as 96, 192, 336, and 720. The best results are bolded and the second best results are underlined.
Models
MSHyper
MSGNet
TimesNet
Crossformer*
DLinear
TiDE*
Pyraformer*
Autoformer
@@
MTGNN
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.383 0.392 0.390 0.411 0.384 0.402 0.418 0.438 0.386 0.400 0.479 0.464 0.664 0.510 0.449 0.459 0.865 0.713 0.440 0.450
192 0.435 0.423 0.442 0.442 0.436 0.429 0.539 0.517 0.437 0.432 0.525 0.492 0.790 0.687 0.500 0.482 1.008 0.792 0.449 0.433
336 0.480 0.445 0.480 0.468 0.491 0.469 0.709 0.638 0.481 0.459 0.565 0.515 0.887 0.738 0.521 0.496 1.107 0.809 0.598 0.554
720 0.482 0.467 0.494 0.488 0.521 0.500 0.733 0.639 0.519 0.516 0.594 0.558 0.976 0.784 0.514 0.512 1.181 0.865 0.685 0.620
ETTh2 96 0.291 0.338 0.328 0.371 0.340 0.374 0.425 0.463 0.333 0.387 0.400 0.440 1.392 0.954 0.346 0.388 3.755 1.525 0.496 0.509
192 0.376 0.391 0.402 0.414 0.402 0.414 0.473 0.500 0.477 0.476 0.528 0.509 3.515 1.561 0.456 0.452 5.602 1.931 0.716 0.616
336 0.419 0.430 0.435 0.443 0.452 0.452 0.581 0.562 0.594 0.541 0.643 0.571 4.471 1.835 0.482 0.486 4.721 1.835 0.718 0.614
720 0.417 0.435 0.417 0.441 0.462 0.468 0.775 0.665 0.831 0.657 0.874 0.679 4.190 1.770 0.515 0.511 3.647 1.625 1.161 0.791
ETTm1 96 0.331 0.350 0.319 0.366 0.338 0.375 0.361 0.403 0.345 0.372 0.364 0.387 0.535 0.510 0.505 0.475 0.672 0.571 0.381 0.415
192 0.374 0.373 0.376 0.397 0.374 0.387 0.387 0.422 0.380 0.389 0.398 0.404 0.580 0.549 0.553 0.496 0.795 0.669 0.442 0.451
336 0.408 0.395 0.417 0.422 0.410 0.411 0.605 0.572 0.413 0.413 0.428 0.425 0.736 0.646 0.621 0.537 1.212 0.871 0.475 0.475
720 0.473 0.431 0.481 0.458 0.473 0.450 0.703 0.645 0.474 0.453 0.487 0.461 1.056 0.782 0.671 0.561 1.166 0.823 0.531 0.507
ETTm2 96 0.179 0.257 0.177 0.262 0.187 0.267 0.275 0.358 0.193 0.292 0.207 0.305 0.361 0.450 0.255 0.339 0.365 0.453 0.240 0.343
192 0.247 0.305 0.247 0.307 0.249 0.309 0.345 0.400 0.284 0.362 0.290 0.364 0.720 0.667 0.281 0.340 0.533 0.563 0.398 0.454
336 0.309 0.344 0.312 0.346 0.321 0.351 0.657 0.528 0.369 0.427 0.377 0.422 1.294 0.882 0.339 0.372 1.363 0.887 0.568 0.555
720 0.401 0.398 0.414 0.403 0.408 0.403 1.208 0.753 0.554 0.522 0.558 0.524 4.147 1.547 0.433 0.432 3.379 1.338 1.072 0.767
Flight 96 0.155 0.260 0.183 0.301 0.237 0.350 0.410 0.449 0.221 0.337 0.247 0.440 0.452 0.508 0.204 0.319 0.333 0.405 0.196 0.316
192 0.155 0.258 0.189 0.306 0.224 0.337 0.503 0.512 0.220 0.336 0.303 0.493 0.458 0.507 0.200 0.314 0.358 0.421 0.272 0.379
336 0.164 0.266 0.206 0.320 0.289 0.394 0.544 0.532 0.229 0.342 0.354 0.533 0.483 0.522 0.201 0.318 0.398 0.446 0.260 0.369
720 0.202 0.299 0.253 0.358 0.310 0.408 0.702 0.635 0.263 0.366 0.572 0.689 0.521 0.536 0.345 0.426 0.476 0.484 0.390 0.449
Weather 96 0.157 0.198 0.163 0.212 0.172 0.220 0.232 0.302 0.196 0.255 0.202 0.261 0.211 0.295 0.266 0.336 0.300 0.384 0.171 0.231
192 0.207 0.244 0.212 0.254 0.219 0.261 0.371 0.410 0.237 0.296 0.242 0.298 0.240 0.316 0.307 0.367 0.598 0.544 0.215 0.274
336 0.265 0.286 0.272 0.299 0.280 0.306 0.495 0.515 0.283 0.335 0.287 0.335 0.295 0.356 0.359 0.395 0.578 0.523 0.266 0.313
720 0.342 0.339 0.350 0.348 0.365 0.359 0.526 0.542 0.345 0.381 0.351 0.386 0.368 0.407 0.419 0.428 1.059 0.741 0.344 0.375
Exchange 96 0.083 0.199 0.102 0.230 0.107 0.234 0.256 0.367 0.088 0.218 0.094 0.218 1.468 1.002 0.197 0.323 0.847 0.752 0.267 0.378
192 0.173 0.294 0.195 0.317 0.226 0.344 0.470 0.509 0.176 0.315 0.184 0.307 1.583 1.061 0.300 0.369 1.204 0.895 0.590 0.578
336 0.310 0.411 0.359 0.436 0.367 0.448 1.268 0.883 0.313 0.427 0.349 0.431 1.733 1.110 0.509 0.524 1.672 1.036 0.939 0.749
720 0.846 0.692 0.940 0.738 0.964 0.746 1.767 1.068 0.839 0.695 0.852 0.698 2.000 1.184 1.447 0.941 2.478 1.310 1.107 0.834
Electricity 96 0.152 0.252 0.165 0.274 0.168 0.272 0.219 0.314 0.197 0.282 0.237 0.329 0.372 0.441 0.201 0.317 0.274 0.368 0.211 0.305
192 0.171 0.271 0.184 0.292 0.184 0.289 0.231 0.322 0.196 0.285 0.236 0.330 0.370 0.440 0.222 0.334 0.296 0.386 0.225 0.319
336 0.187 0.284 0.195 0.302 0.198 0.300 0.246 0.337 0.209 0.301 0.249 0.344 0.368 0.440 0.231 0.338 0.300 0.394 0.247 0.340
720 0.224 0.316 0.231 0.332 0.220 0.320 0.280 0.363 0.245 0.333 0.284 0.373 0.373 0.446 0.254 0.361 0.373 0.439 0.287 0.373
  • * indicates that some methods do not have uniform prediction lengths with other methods. To ensure a fair comparison, we utilize their official code and adjust prediction lengths. Other results are from MSGNet.
  • MSHyper gets the SOTA results on all eight datasets and even achieves better performance on the datasets with low auto-correlation function (ACF) values (e.g., Exchange-Rate and ETTm1 datasets). Specifically, MSHyper reduces prediction errors by an average of 4.06% and 5.27% over the best baseline in MSE and MAE, respectively. We attribute this to the reason that MSHyper can get multi-scale embeddings and capture high-order interactions between temporal patterns of different scales. In addition, prediction errors increase smoothly and slowly when the forecasting horizon is increased, which means that the MSHyper retains better long-range robustness. These observations provide empirical guidance for the success of using MSHyper in multivariate long-range time series forecasting.
  • Traditional transformer-based methods (i.e., Informer, Autoformer, and Pyraformer) exhibit relatively poor predictive performance. This is because time series forecasting requires modeling temporal pattern interactions, while vanilla attention or simplistic decomposition techniques are insufficient in capturing such interactions.
  • Although using different backbones, existing multi-scale decomposition methods (i.e., MTGNN, DLinear, TiDE, Crossformer, TimesNet, and MSGNet) all achieve competitive performance. Specially, by considering multi-scale pattern interactions on the decomposed sequences, MSGNet and TimesNet achieve promising forecasting results. However, they fail to consider high-order pattern interactions and get worse performance than MSHyper in most cases.
Univariate Results. Table IV summarizes the univariate long-range time series forecasting results of MSHyper compared with baselines on ETT dataset, from which we can discern the following tendencies:
  • MSHyper still achieves the SOTA performance for univariate long-range time series forecasting, and the prediction errors increase steadily and slowly when increasing the forecasting horizon, which demonstrates the effectiveness of MSHyper in improving the capability of univariate long-range time series forecasting.
  • Although MLP-based methods perform better than traditional Transformer-based methods, they fail to capture multi-scale pattern interactions and get worse performance than MSGNet, TimesNet, and MSHyper.
  • MSGNet and TimesNet are the best baselines that use multi-head attention mechanism and 2D convolution kernels to learn multi-scale pattern interactions, respectively. However, they only consider one type of multi-scale (i.e., intra-scale or mixed-scale) pattern interactions and get worse performance than MSHyper. In contrast, MSHyper leverages multi-scale hypergraph structures to capture intra-scale, inter-scale, and mixed-scale pattern interactions.
TABLE IV: Univariate long-range time series forecasting results on ETT dataset. The input length is set as $I$=96, and the prediction length $O$ is set as 96, 192, 336, and 720. The best results are bolded and the second best results are underlined.
Models
MSHyper
MSGNet
TimesNet
Crossformer
DLinear
TiDE
Pyraformer
@@
Informer
MTGNN
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.055 0.178 0.059 0.186 0.058 0.185 0.076 0.216 0.061 0.184 0.068 0.193 0.461 0.621 0.449 0.459 0.865 0.713 0.382 0.560
192 0.076 0.209 0.078 0.214 0.077 0.213 0.085 0.225 0.083 0.214 0.083 0.216 0.620 0.731 0.500 0.482 1.008 0.792 0.470 0.624
336 0.090 0.237 0.096 0.242 0.104 0.249 0.106 0.257 0.100 0.245 0.098 0.244 0.494 0.650 0.521 0.496 1.107 0.809 0.381 0.553
720 0.098 0.246 0.126 0.278 0.086 0.232 0.128 0.287 0.197 0.369 0.178 0.347 0.697 0.784 0.514 0.512 1.181 0.865 0.313 0.498
ETTh2 96 0.127 0.274 0.144 0.295 0.141 0.293 0.138 0.289 0.128 0.271 0.145 0.296 0.405 0.519 0.358 0.397 3.755 1.525 0.516 0.620
192 0.178 0.327 0.198 0.353 0.192 0.345 0.188 0.341 0.182 0.328 0.179 0.346 1.528 1.102 0.456 0.452 5.602 1.931 1.021 0.902
336 0.217 0.372 0.224 0.381 0.240 0.394 0.229 0.384 0.231 0.377 0.226 0.367 1.546 1.110 0.482 0.486 4.721 1.835 1.496 1.108
720 0.228 0.383 0.222 0.380 0.238 0.390 0.250 0.404 0.322 0.464 0.319 0.462 1.302 1.016 0.515 0.511 3.647 1.625 1.969 1.306
ETTm1 96 0.028 0.125 0.028 0.126 0.029 0.127 0.071 0.193 0.034 0.135 0.034 0.136 0.089 0.239 0.505 0.475 0.672 0.571 0.112 0.271
192 0.042 0.156 0.044 0.160 0.047 0.163 0.106 0.245 0.054 0.174 0.055 0.174 0.214 0.388 0.553 0.496 0.795 0.669 0.217 0.393
336 0.056 0.181 0.059 0.187 0.058 0.187 0.104 0.287 0.075 0.204 0.073 0.202 0.332 0.505 0.621 0.537 1.212 0.871 0.389 0.556
720 0.078 0.215 0.081 0.218 0.081 0.220 0.183 0.334 0.113 0.254 0.100 0.236 0.543 0.662 0.671 0.561 1.166 0.823 0.633 0.719
ETTm2 96 0.069 0.189 0.074 0.206 0.075 0.201 0.071 0.194 0.070 0.191 0.073 0.197 0.167 0.314 0.255 0.339 0.365 0.453 0.156 0.313
192 0.099 0.235 0.108 0.249 0.107 0.248 0.105 0.242 0.104 0.238 0.107 0.242 0.550 0.603 0.281 0.340 0.533 0.563 0.194 0.341
336 0.127 0.270 0.140 0.287 0.142 0.289 0.135 0.281 0.134 0.277 0.138 0.282 0.873 0.757 0.339 0.372 1.363 0.887 0.816 0.809
720 0.176 0.326 0.194 0.343 0.206 0.352 0.189 0.340 0.186 0.331 0.188 0.333 2.449 1.400 0.422 0.419 3.379 1.338 1.193 0.884
V-D Ablation Studies
Multi-Scale Feature Extraction. Our multi-scale feature extraction is implemented by 1D convolution. To investigate the effect of the multi-scale feature extraction, we conduct ablation studies by carefully designing the following two variants:
  • MSHyper-avg: It replaces 1D convolution with avg-pooling to extract multi-scale features.
  • MSHyper-max: It replaces 1D convolution with max-pooling to extract multi-scale features.
The experimental results are shown in Table V, from which we can observe that MSHyper achieves the best performance in almost all cases, which indicates that convolution kernels can extract more complex features.
TABLE V: Results of different multi-scale feature extraction methods.
Methods Predicition Length
96 192 336 720
MSHyper-max
MSE
MAE
0.384
0.393
0.437
0.426
0.483
0.446
0.487
0.464
MSHyper-avg
MSE
MAE
0.382
0.393
0.436
0.426
0.478
0.447
0.485
0.468
@@ V-D Ablation Studies
MSE
MAE
0.383
0.392
0.435
0.423
0.480
0.445
0.482
0.467
Multi-Scale Hypergraph. To investigate the effectiveness of the multi-scale hypergraph, we conduct ablation studies by carefully designing the following three variants:
  • MSHyper-FCG: It replaces the H-HGC model with the fully-connected graph.
  • MSHyper-PG: It replaces the H-HGC model with the pyramid graph used in Pyraformer [28].
  • MSHyper-SSH: It connects the input sequence with intra-scale hyperedges, and thus the multi-scale hypergraph turns to the single-scale hypergraph.
The experimental results are shown in Table VI, from which we can observe that MSHyper performs the best in all cases, which indicates the importance of multi-scale hypergraph in modeling high-order interactions between temporal patterns of different scales.
TABLE VI: Results of different multi-scale hypergraph construction methods.
Methods Predicition Length
96 192 336 720
@@
0.383
0.392
0.435
0.423
0.480
0.445
0.482
0.467
Hyperedge Graph. To investigate the effectiveness of the hyperedge graph, we conduct ablation studies by carefully designing the following three variants:
  • MSHyper-w/o AR: It removes the association relationship of the hyperedge graph.
  • MSHyper-w/o SR: It removes the sequential relationship of the hyperedge graph.
  • MSHyper-w/o H-H: It removes the hyperedge graph (i.e., without the hyperedge-hyperedge phase).
The experimental results are shown in Table VII, from which we can discern the following tendencies: 1) MSHyper performs better than MSHyper-w/o AR and MSHyper-w/o SR, showing the effectiveness of the association relationship and sequential relationship, respectively. 2) Removing the hyperedge-hyperedge phase gets the worst forecasting results, which demonstrates the superiority of the hyperedge graph in enhancing hypergraph modeling.
TABLE VII: Results of different hyperedge graph construction methods.
Methods Predicition Length
96 192 336 720